OpenStudy (anonymous):

I need urgent help with this exponential function 24p^3q x 3p^5q^-3 (divide) 27p^-6q^4 PLEASE HELP!!! TEST TOMORROW

OpenStudy (anonymous):

8p^6q^4+8 /3q^2

OpenStudy (anonymous):

the answer says 8p^14 / 3q^6 What do you mean in latex?

OpenStudy (campbell_st):

it looks like \[\frac{24 p^3 q \times 3p^3q^{-3}}{27p^{-6}q^4}\] so use number, letters, powers, signs

OpenStudy (anonymous):

oh ok.. thank you

OpenStudy (campbell_st):

so the numbers \[\frac{24 \times 3}{27} = \frac{72}{27} = \frac{8}{3}\] next the letters \[\frac{8pq}{3}\] now the signs when multiplying add the powers when dividing subtract the powers

OpenStudy (campbell_st):

power of p 3 + 5 - (-3) = 11 powers of q 1 + -3 - 4 = -6 so its \[\frac{8p^{11}q^{-6}}{3}\] last step.. negative powers indicate fractions so its \[q^{-6} = \frac{1}{q^6}\] so the answer becomes \[\frac{8p^{11}}{3q^6}\] hope it makes sense

OpenStudy (anonymous):

8p^14 q^2 /3

OpenStudy (anonymous):

Thanks very much.

OpenStudy (anonymous):

do you guys know much about exponential functions?

OpenStudy (anonymous):

No I don't, just learning as I go.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I just have questions about them that's all